OpenStudy (anonymous):

Jim has invested $10000 at a rate of 7.2% that is compounded monthly. How many years will it take for this amount to grow to $25000?

OpenStudy (anonymous):

A=P(1+r/n)^(nt) \[25000=10000(1+\frac{ .072 }{ 12 })^{12t}\]